The brick wall (of thermal conductivity
0.89 W/m ·◦ C) of a building has dimensions
of 2.6 m by 6 m and is 22 cm thick.
How much heat flows through the wall in
a 11.1 h period when the average inside and
outside temperatures are, respectively, 20◦C
and 8◦C?
Answer in units of MJ.
1
Expert's answer
2020-04-06T09:17:48-0400
"E=Pt=\frac{kA\Delta T}{L}t"

"E=\frac{(0.89)(2.6)(6)(20-8)}{0.22}(11.1\cdot 3600)=30\cdot 10^6 J"

"E=30\ MJ"